Determine where the radial nodes in a 3s orbital of the hydrogen atom would be based on the radial wave equation. Recall that the number of nodes is equal to n-1.
Answer #1

For 3s orbital : n = 3 , L= 0 ,

for radial node of 3s orbital = n - L-1 = 3- 0 - 1 = 2

total nodes = angular node + radial node = L + n - L - 1 = n - 1 = 3 - 1 = 2